\begin{thebibliography}{99}
%-------------------------------------------------------------------------------------------------------------------------
%
%[1]
\bibitem{Bierenbaum:2009mv}
  I.~Bierenbaum, J.~Bl\"umlein and S.~Klein,
  %``Mellin Moments of the O(alpha**3(s)) Heavy Flavor Contributions to unpolarized Deep-Inelastic Scattering at Q**2 >> m**2 and Anomalous Dimensions,''
  Nucl.\ Phys.\ B {\bf 820} (2009) 417
  [arXiv:0904.3563 [hep-ph]].
  %%CITATION = ARXIV:0904.3563;%%
%-------------------------------------------------------------------------------------------------------------------------
%
%[2]
\bibitem{Behring:2014eya}
  A.~Behring, I.~Bierenbaum, J.~Bl\"umlein, A.~De Freitas, S.~Klein and F.~Wi\ss{}brock,
  %``The logarithmic contributions to the $O(\alpha^3_s)$ asymptotic massive Wilson coefficients and operator matrix elements in deeply inelastic scattering,''
  Eur.\ Phys.\ J.\ C {\bf 74} (2014) 9,  3033
  [arXiv:1403.6356 [hep-ph]].
  %%CITATION = ARXIV:1403.6356;%%
%-------------------------------------------------------------------------------------------------------------------------
%
%[3]
\bibitem{Ablinger:2010ty}
  J.~Ablinger, J.~Bl\"umlein, S.~Klein, C.~Schneider and F.~Wi\ss{}brock,
  %``The O(\alpha_s^3) Massive Operator Matrix Elements of O(n_f) for the Structure Function F_2(x,Q^2) and Transversity,''
  Nucl.\ Phys.\ B {\bf 844} (2011) 26
  [arXiv:1008.3347 [hep-ph]].
  %%CITATION = ARXIV:1008.3347;%%
%------------------------------------------------------------------------
%
%[4]
\bibitem{Ablinger:2014vwa}
  J.~Ablinger, A.~Behring, J.~Bl\"umlein, A.~De Freitas, A.~Hasselhuhn, A.~von Manteuffel, 
  M.~Round, C.~Schneider, and F.~Wi\ss{}brock,
  %``The 3-Loop Non-Singlet Heavy Flavor Contributions and Anomalous Dimensions for the Structure Function $F_2(x,Q^2)$ and Transversity,''
  Nucl.\ Phys.\ B {\bf 886} (2014) 733
  [arXiv:1406.4654 [hep-ph]];\\
  %%CITATION = ARXIV:1406.4654;%%
%\bibitem{Behring:2015zaa}
  A.~Behring, J.~Bl\"umlein, A.~De Freitas, A.~von Manteuffel and C.~Schneider,
  %``The 3-Loop Non-Singlet Heavy Flavor Contributions to the Structure Function $g_{1}(x,Q^{2})$ at Large Momentum Transfer,''
  Nucl.\ Phys.\ B {\bf 897} (2015) 612
%  doi:10.1016/j.nuclphysb.2015.06.007
  [arXiv:1504.08217 [hep-ph]].
  %%CITATION = doi:10.1016/j.nuclphysb.2015.06.007;%%
%------------------------------------------------------------------------
%
%[5]
\bibitem{Ablinger:2014nga}
  J.~Ablinger, A.~Behring, J.~Bl\"umlein, A.~De Freitas, A.~von Manteuffel and C.~Schneider,
  %``The 3-loop pure singlet heavy flavor contributions to the structure function $F_2(x,Q^2)$ and the anomalous dimension,''
  Nucl.\ Phys.\ B {\bf 890} (2014) 48
  [arXiv:1409.1135 [hep-ph]].
  %%CITATION = ARXIV:1409.1135;%%
%------------------------------------------------------------------------
%
%[6]
\bibitem{Ablinger:2014lka}
  J.~Ablinger, J.~Bl\"umlein, A.~De Freitas, A.~Hasselhuhn, A.~von Manteuffel, M.~Round, C.~Schneider and 
  F.~Wi\ss{}brock,
  %``The Transition Matrix Element A_{gq}(N) of the Variable Flavor Number Scheme at O(\alpha_s^3),''
  Nucl.\ Phys.\ B {\bf 882} (2014) 263
  [arXiv:1402.0359 [hep-ph]].
  %%CITATION = ARXIV:1402.0359;%%
%------------------------------------------------------------------------
%
%[7]
\bibitem{Ablinger:2016swq}
  J.~Ablinger {\it et al.},
  %``New Results on Massive 3-Loop Wilson Coefficients in Deep-Inelastic Scattering,''
  PoS LL {\bf 2016} (2016) 065
  [arXiv:1609.03397 [hep-ph]].
  %%CITATION = ARXIV:1609.03397;%%
%-------------------------------------------------------------------------------------------------------------------------
%
%[8]
\bibitem{Alekhin:2012vu}
  S.~Alekhin, J.~Bl\"umlein, K.~Daum, K.~Lipka and S.~Moch,
  %``Precise charm-quark mass from deep-inelastic scattering,''
  Phys.\ Lett.\ B {\bf 720} (2013) 172
%  doi:10.1016/j.physletb.2013.02.010
  [arXiv:1212.2355 [hep-ph]].
  %%CITATION = doi:10.1016/j.physletb.2013.02.010;%%
%-------------------------------------------------------------------------------------------------------------------------
%
%[9]
\bibitem{Alekhin:2017kpj}
  S.~Alekhin, J.~Bl\"umlein, S.~Moch and R.~Placakyte,
  %``Parton distribution functions, $\alpha_s$, and heavy-quark masses for LHC Run II,''
  Phys.\ Rev.\ D {\bf 96} (2017) no.1,  014011
%  doi:10.1103/PhysRevD.96.014011
  [arXiv:1701.05838 [hep-ph]].
  %%CITATION = doi:10.1103/PhysRevD.96.014011;%%
%-------------------------------------------------------------------------------------------------------------------------
%
%[10]
\bibitem{Ablinger:2017err}
  J.~Ablinger, J.~Bl\"umlein, A.~De Freitas, A.~Hasselhuhn, C.~Schneider and F.~Wi\ss{}brock,
  %``Three Loop Massive Operator Matrix Elements and Asymptotic Wilson Coefficients with Two Different Masses,''
  Nucl.\ Phys.\ B {\bf 921} (2017) 585
%  doi:10.1016/j.nuclphysb.2017.05.017
  [arXiv:1705.07030 [hep-ph]].
  %%CITATION = doi:10.1016/j.nuclphysb.2017.05.017;%%
%-------------------------------------------------------------------------------------------------------------------------
%
%[11]
\bibitem{Ablinger:2011pb}
  J.~Ablinger, J.~Bl\"umlein, S.~Klein, C.~Schneider and F.~Wi\ss{}brock,
  %``3-Loop Heavy Flavor Corrections to DIS with two Massive Fermion Lines,''
  arXiv:1106.5937 [hep-ph].
  %%CITATION = ARXIV:1106.5937;%%
%------------------------------------------------------------------------
%
%[12]
\bibitem{Ablinger:2012qj}
  J.~Ablinger, J.~Bl\"umlein, A.~Hasselhuhn, S.~Klein, C.~Schneider and F.~Wi\ss{}brock,
  %``New Heavy Flavor Contributions to the DIS Structure Function $F_2(x,Q^2)$ at $O(\alpha_s^3),''
  PoS {(RADCOR2011)} 031
  [arXiv:1202.2700 [hep-ph]].
%------------------------------------------------------------------------
%
%[13]
\bibitem{Buza:1995ie}
  M.~Buza, Y.~Matiounine, J.~Smith, R.~Migneron and W.~L.~van Neerven,
  %``Heavy quark coefficient functions at asymptotic values Q**2 >> m**2,''
  Nucl.\ Phys.\ B {\bf 472} (1996) 611
%  doi:10.1016/0550-3213(96)00228-3
  [hep-ph/9601302].
  %%CITATION = doi:10.1016/0550-3213(96)00228-3;%%
%-------------------------------------------------------------------------------------------------------------------------
%
%[14]
\bibitem{CS1}
%\bibitem{Chetyrkin:1999qi}
  K.G.~Chetyrkin and M.~Steinhauser,
  %``The relation between the MS-bar and the on-shell quark mass at order
  %alpha(s)**3,''
  Nucl.\ Phys.\  B {\bf 573}, 617 (2000)
  [arXiv:hep-ph/9911434];
  %%CITATION = NUPHA,B573,617;%%
%\bibitem{Chetyrkin:1999ys}
%  K.G.~Chetyrkin and M.~Steinhauser,
  %``Short distance mass of a heavy quark at order alpha(s)**3,''
  Phys.\ Rev.\ Lett.\  {\bf 83}, 4001 (1999)  
  [arXiv:hep-ph/9907509].
  %%CITATION = PRLTA,83,4001;%% 
%------------------------------------------------------------------------
%
%[15]
\bibitem{YND}
F.J.~Yndurain, {\sf The Theory of Quark and Gluon Interactions}, (Springer, Berlin, 1983).
%------------------------------------------------------------------------
%
%[16]
\bibitem{Kallen:1955fb}
  A.O.G.~K\"allen and A.~Sabry,
  %``Fourth order vacuum polarization,''
  Kong.\ Dan.\ Vid.\ Sel.\ Mat.\ Fys.\ Med.\  {\bf 29} (1955) no.17,  1.
%  doi:10.1007/978-3-319-00627-7_93
  %%CITATION = doi:10.1007/978-3-319-00627-7_93;%%
%------------------------------------------------------------------------
%
%[17]
\bibitem{Broadhurst:1991fi}
  D.~J.~Broadhurst,
  %``Three loop on-shell charge renormalization without integration: Lambda-MS (QED) to four loops,''
%  Z.\ Phys.\ C {\bf 54} (1992) 599.
  doi:10.1007/BF01559486
  %%CITATION = doi:10.1007/BF01559486;%%
%------------------------------------------------------------------------
%
%[18]
\bibitem{Fleischer:1992re}
  J.~Fleischer and O.V.~Tarasov,
  %``Gauge invariant on-shell Z(1) in QED and QCD,''
  Phys.\ Lett.\ B {\bf 283} (1992) 129.
%  doi:10.1016/0370-2693(92)91442-C
  %%CITATION = doi:10.1016/0370-2693(92)91442-C;%%
%------------------------------------------------------------------------
%
%[19]
\bibitem{Jegerlehner:1998zg}
  F.~Jegerlehner and O.V.~Tarasov,
  %``Exact mass dependent two loop anti-alpha(s)(Q**2) in the background MOM renormalization scheme,''
  Nucl.\ Phys.\ B {\bf 549} (1999) 481
%  doi:10.1016/S0550-3213(99)00141-8
  [hep-ph/9809485].
  %%CITATION = doi:10.1016/S0550-3213(99)00141-8;%%
%------------------------------------------------------------------------
%
%[20]
\bibitem{Chetyrkin:2008jk}
  K.G.~Chetyrkin, B.A.~Kniehl and M.~Steinhauser,
  %``Strong-Coupling Constant at Three Loops in Momentum Subtraction Scheme,''
  Nucl.\ Phys.\ B {\bf 814} (2009) 231
%  doi:10.1016/j.nuclphysb.2009.01.026
  [arXiv:0812.1337 [hep-ph]].
  %%CITATION = doi:10.1016/j.nuclphysb.2009.01.026;%%
%------------------------------------------------------------------------
%
%[21]
\bibitem{Nogueira:1991ex}
  P.~Nogueira,
  %``Automatic Feynman graph generation,''
  J.\ Comput.\ Phys.\  {\bf 105} (1993) 279.
  %%CITATION = JCTPA,105,279;%%
%------------------------------------------------------------------------
%
%[22]
\bibitem{vanRitbergen:1998pn}
  T.~van Ritbergen, A.N.~Schellekens and J.A.M.~Vermaseren,
  %``Group theory factors for Feynman diagrams,''
  Int.\ J.\ Mod.\ Phys.\  A {\bf 14}  (1999) 41
  [arXiv:hep-ph/9802376].
  %%CITATION = IMPAE,A14,41;%%
%------------------------------------------------------------------------
%
%[23]
\bibitem{Ablinger:2015tua}
  J.~Ablinger, A.~Behring, J.~Bl\"umlein, A.~De Freitas, A.~von Manteuffel and C.~Schneider,
  %``Calculating Three Loop Ladder and V-Topologies for Massive Operator Matrix Elements by Computer Algebra,''
  Comput.\ Phys.\ Commun.\  {\bf 202} (2016) 33
%  doi:10.1016/j.cpc.2016.01.002
  [arXiv:1509.08324 [hep-ph]].
  %%CITATION = doi:10.1016/j.cpc.2016.01.002;%%
%------------------------------------------------------------------------
%
%[24]
\bibitem{MB1a}
E.W.~Barnes, 
%"A new development of the theory of the hypergeometric functions",
Proc. Lond. Math. Soc. (2) {\bf 6} (1908) 141.
%     pages     = "141-177",
%------------------------------------------------------------------------
%
%[25]
\bibitem{MB1b}
E.W.~Barnes,
%     title     = "A transformation of generalised hypergeometric series",
Quarterly Journal of Mathematics {\bf 41} (1910) 136.
%     pages     = "136-140"
%------------------------------------------------------------------------
%
%[26]
\bibitem{MB2}
H.~Mellin,
%     title     = "Abri{\ss} einer einheitlichen Theorie der Gamma- und der hypergeometrischen Funktionen",
Math. Ann. {\bf 68}, no. 3 (1910) 305.
%     pages     = "305???337",
%------------------------------------------------------------------------
%
%[27]
\bibitem{MB3}
E.T.~Whittaker and G.N.~Watson, {\sf A Course of Modern Analysis}, (Cambridge University Press, Cambridge, 1927;
                   reprinted 1996) 616~p.
%------------------------------------------------------------------------
%
%[28]
\bibitem{MB4}
E.C.~Titchmarsh,
{\sf Introduction to the Theory of Fourier Integrals},
(Calendron Press, Oxford, 1937; 2nd Edition 1948).
%------------------------------------------------------------------------
%
%[29]
\bibitem{Czakon:2005rk}
  M.~Czakon,
  %``Automatized analytic continuation of Mellin-Barnes integrals,''
  Comput.\ Phys.\ Commun.\  {\bf 175} (2006) 559
  [hep-ph/0511200];\\
  %%CITATION = HEP-PH/0511200;%%
%\bibitem{Smirnov:2009up}
  A.V.~Smirnov and V.A.~Smirnov,
  %``On the Resolution of Singularities of Multiple Mellin-Barnes Integrals,''
  Eur.\ Phys.\ J.\ C {\bf 62} (2009) 445
  [arXiv:0901.0386 [hep-ph]].
  %%CITATION = ARXIV:0901.0386;%%
%------------------------------------------------------------------------
%
%[30]
\bibitem{SIG1}
C.~Schneider, {S\'em.~Lothar. Combin.\/} {\bf 56} (2007) 1, %--36
 article B56b.
%-----------------------------------------------------------------------------------
%
%[31]
\bibitem{SIG2}
C.~Schneider, {{\sf Computer Algebra in Quantum Field Theory: Integration,
  Summation and Special Functions}\/} Texts and Monographs in Symbolic
  Computation eds. C.~Schneider and J.~Bl\"umlein  (Springer, Wien, 2013) 325, %--360
  arXiv:1304.4134 [cs.SC].
%-----------------------------------------------------------------------------------
%
%[32]
\bibitem{EMSSP}
%\bibitem{Ablinger:2010pb}
  J.~Ablinger, J.~Bl\"umlein, S.~Klein and C.~Schneider,
  %``Modern Summation Methods and the Computation of 2- and 3-loop Feynman Diagrams,''
  Nucl.\ Phys.\ Proc.\ Suppl.\  {\bf 205-206} (2010) 110
  [arXiv:1006.4797 [math-ph]];\\
  %%CITATION = ARXIV:1006.4797;%%
%\bibitem{Blumlein:2012hg}
  J.~Bl\"umlein, A.~Hasselhuhn and C.~Schneider,
  %``Evaluation of Multi-Sums for Large Scale Problems,''
  PoS (RADCOR 2011) 032
  [arXiv:1202.4303 [math-ph]];\\
  %%CITATION = ARXIV:1202.4303;%%
  C. Schneider,  % Symbolic Summation in Difference Fields and Its Application in Particle Physics.
  Computer Algebra Rundbrief {\bf 53} (2013), 8%-12
  ;\\
  %\bibitem{Schneider:2013zna}
  C.~Schneider,
  %``Modern Summation Methods for Loop Integrals in Quantum Field Theory: The Packages Sigma, EvaluateMultiSums and SumProduction,''
  J.\ Phys.\ Conf.\ Ser.\  {\bf 523} (2014) 012037
  [arXiv:1310.0160 [cs.SC]].
  %%CITATION = ARXIV:1310.0160;%%
%-----------------------------------------------------------------------------------
%
%[33]
\bibitem{HARMONICSUMS}
%\bibitem{Ablinger:2014dda}
  J.~Ablinger,
  %``The package HarmonicSums: Computer Algebra and Analytic Aspects of Nested Sums,''
  PoS  (LL2014) 019;
  %%CITATION = POSCI,LL2014,019;%%
%\bibitem{Ablinger:2010kw}
%  J.~Ablinger,
  {\sf A Computer Algebra Toolbox for Harmonic Sums Related to Particle Physics}, Diploma Thesis,
J. Kepler University Linz, 2009,
  arXiv:1011.1176 [math-ph].
  %%CITATION = ARXIV:1011.1176;%%
%--------------------------------------------------------------------------------------------------
%
%[34]
\bibitem{Ablinger:PhDThesis}
  J.~Ablinger,
  {\it Computer Algebra Algorithms for Special Functions in Particle Physics},
  Ph.D. Thesis, J. Kepler University Linz, 2012,
  arXiv:1305.0687 [math-ph];
  %%CITATION = ARXIV:1305.0687;%%
%--------------------------------------------------------------------------------------------------
%
%[35]
\bibitem{Ablinger:2013cf}
  J.~Ablinger, J.~Bl\"umlein and C.~Schneider,
  %``Analytic and Algorithmic Aspects of Generalized Harmonic Sums and Polylogarithms,''
  J.\ Math.\ Phys.\  {\bf 54} (2013) 082301
  [arXiv:1302.0378 [math-ph]].
  %%CITATION = ARXIV:1302.0378;%%
%--------------------------------------------------------------------------------------------------
%
%[36]
\bibitem{Ablinger:2011te}
  J.~Ablinger, J.~Bl\"umlein and C.~Schneider,
  %``Harmonic Sums and Polylogarithms Generated by Cyclotomic Polynomials,''
  J.\ Math.\ Phys.\  {\bf 52} (2011) 102301
  [arXiv:1105.6063 [math-ph]].
  %%CITATION = ARXIV:1105.6063;%%
%--------------------------------------------------------------------------------------------------
%
%[37]
\bibitem{Ablinger:2014bra}
  J.~Ablinger, J.~Bl\"umlein, C.G.~Raab and C.~Schneider,
  %``Iterated Binomial Sums and their Associated Iterated Integrals,''
  J.\ Math.\ Phys.\  {\bf 55} (2014) 112301
%  doi:10.1063/1.4900836
  [arXiv:1407.1822 [hep-th]].
  %%CITATION = doi:10.1063/1.4900836;%%
%----------------------------------------------------------------------------
%
%[38]
\bibitem{Steinhauser:2000ry}
    M.~Steinhauser,
    %``MATAD: A program package for the computation of massive tadpoles,''
    Comput.\ Phys.\ Commun.\  {\bf 134} (2001) 335,
    [arXiv:hep-ph/0009029]; code~{\sf MATAD 3.0}.
    %%CITATION = CPHCB,134,335;%%
%----------------------------------------------------------------------------
%
%[39]
\bibitem{Harlander:1997zb}
  R.~Harlander, T.~Seidensticker and M.~Steinhauser,
  %``Complete corrections of Order alpha alpha-s to the decay of the Z boson into bottom quarks,''
  Phys.\ Lett.\ B {\bf 426} (1998) 125
  [hep-ph/9712228].
  %%CITATION = HEP-PH/9712228;%%
%------------------------------------------------------------------------
%
%[40]
\bibitem{Seidensticker:1999bb}
  T.~Seidensticker,
  %``Automatic application of successive asymptotic expansions of Feynman diagrams,''
  hep-ph/9905298.
  %%CITATION = HEP-PH/9905298;%%
%------------------------------------------------------------------------
%
%[41]
\bibitem{Lee:2013sx}
  R.~Lee, P.~Marquard, A.V.~Smirnov, V.A.~Smirnov and M.~Steinhauser,
  %``Four-loop corrections with two closed fermion loops to fermion self energies and the lepton anomalous magnetic moment,''
  JHEP {\bf 1303} (2013) 162
%  doi:10.1007/JHEP03(2013)162
  [arXiv:1301.6481 [hep-ph]].
  %%CITATION = doi:10.1007/JHEP03(2013)162;%%
%------------------------------------------------------------------------
\end{thebibliography}

\documentclass[12pt]{article}  
\usepackage{cite}
\usepackage{epsfig}
\usepackage{graphicx}
\usepackage{amsmath}
\usepackage{amssymb}
\usepackage{ulem}
\usepackage{mdwlist}          
\usepackage{color}              

\usepackage{a41}
\usepackage{color}
\usepackage[rflt]{floatflt}
\usepackage{float}
\usepackage{slashed}

\newcommand{\lnMa}[1]{\ln^{#1}\biggl(\frac{m_1^2}{\mu^2}\biggr)}
\newcommand{\lnMb}[1]{\ln^{#1}\biggl(\frac{m_2^2}{\mu^2}\biggr)}

\usepackage{array}

\usepackage[english]{babel}
\usepackage[latin1]{inputenc}
\usepackage[T1]{fontenc}
\usepackage{ae}

\usepackage{url}

\usepackage{amsmath, amsthm, amssymb}
\newtheorem{thm}{Theorem}[section]

\newtheorem{definition}[thm]{Definition}

\usepackage{rotating}

\usepackage{graphicx}

\newcounter{mmacnt}
\def\restartmma{\setcounter{mmacnt}{0}}
\restartmma \catcode`|=\active
\def|#1|{\mathrm{#1}}
\catcode`|=12
\newenvironment{mma}{
 \par\smallskip
 \catcode`|=\active
 \parskip=0pt\parindent=0pt 
 \small
 \def\In##1\\{%
   \def\linebreak{\hfill\break\null\qquad}%
   \refstepcounter{mmacnt}
   \hangindent=2.5em\hangafter=0
   \leavevmode
   \llap{\tiny\sffamily In[\arabic{mmacnt}]:=\kern.5em}%
   \mathversion{bold}\footnotesize$\displaystyle##1$\normalsize
   \mathversion{normal}\par
 }%
 \def\Print##1\\{%
   \def\linebreak{\hfill\break}%
   \hangindent=2.5em\hangafter=0
   \leavevmode ##1\par}%
 \def\Out##1\\{%
   \def\linebreak{$\hfill\break\null\hfill$}%
   \kern\abovedisplayskip\par
   \hangindent=2.5em\hangafter=0
   \leavevmode
   \llap{\tiny\sffamily Out[\arabic{mmacnt}]=\kern.5em}
   \footnotesize$\displaystyle##1$\normalsize\hfill\null\par
   \kern\belowdisplayskip
 }%
 \def\Warning##1##2\\{%
   \def\linebreak{\hfill\break}%
   \hangindent=2.5em\hangafter=0
   \leavevmode
   {\scriptsize##1 : ##2}\par}%
}{%
 \par\smallskip
}

\usepackage{color}

\newenvironment{fshaded}{%
\MakeFramed {\FrameRestore}
}%
{\endMakeFramed}

\allowdisplaybreaks[4]

\begin{document}
\setlength{\baselineskip}{0.515cm}
\sloppy
\thispagestyle{empty}
\begin{flushleft}
DESY 17--157
\\
DO--TH 17/26\\
October  2017\\
\end{flushleft}

\mbox{}
\vspace*{\fill}
\begin{center}

{\LARGE\bf The Three Loop Two-Mass Contribution to the} 

\vspace*{3mm} 
{\LARGE\bf Gluon Vacuum Polarization}

\vspace{3cm}
\large
J.~Bl\"umlein$^a$, 
A.~De Freitas$^a$, 
C.~Schneider$^b$,
K.~Sch\"onwald$^a$

\vspace{1.cm}
\normalsize
{\it  $^a$ Deutsches Elektronen--Synchrotron, DESY,}\\
{\it  Platanenallee 6, D-15738 Zeuthen, Germany}
\\

\vspace*{3mm}
{\it $^b$~Research Institute for Symbolic Computation (RISC),\\
                          Johannes Kepler University, Altenbergerstra\ss{}e 69,
                          A--4040, Linz, Austria}\\


\end{center}
\normalsize
\vspace{\fill}
\begin{abstract}
\noindent
We calculate the two-mass contribution to the 3-loop vacuum polarization of the gluon in 
Quantum Chromodynamics at virtuality $p^2 = 0$ for general masses and also present the 
analogous result for the photon in Quantum Electrodynamics. 
\end{abstract}

\vspace*{\fill}
\noindent

\newpage
The 3-loop heavy flavor corrections to deep-inelastic scattering at larger virtualities \cite{
Bierenbaum:2009mv,Behring:2014eya,Ablinger:2010ty,Ablinger:2014vwa,Ablinger:2014nga,
Ablinger:2014lka,Ablinger:2016swq} form an important ingredient for the determination of the strong 
coupling constant $\alpha_s(M_Z^2)$, the parton distribution functions and the measurement of the mass 
of the charm quark $m_c$ at high precision \cite{Alekhin:2012vu,Alekhin:2017kpj}. Starting at 3-loop 
order the QCD corrections contain also 2-mass contributions in single Feynman diagrams 
\cite{Ablinger:2017err,Ablinger:2011pb,Ablinger:2012qj}. The heavy flavor contributions 
to deep-inelastic structure functions in the region of larger virtualities $Q^2 \gg m_Q^2$, with $m_Q$ the 
heavy quark mass, can be obtained in terms of massive on-shell operator matrix elements (OMEs) 
\cite{Buza:1995ie,Bierenbaum:2009mv}. These quantities also receive massive self-energy insertions, such as the
on-shell vacuum polarization function
$\hat{\tilde{\Pi}}^{(3)}(0,m_1^2,m_2^2,\mu^2)$ and fermion self energy $\hat{\tilde{\Sigma}}(0,m_1^2,m_2^2,
\mu^2)$. Here $m_{1,2}$ denote the corresponding heavy quark masses and $\mu$ is the renormalization scale.
These quantities are of more general interest, as they appear in various massive higher loop calculations.
The expression for  $\hat{\tilde{\Sigma}}(0,m_1^2,m_2^2,\mu^2)$ for general ratios $\eta = m_1^2/m_2^2$
has been given in Ref.~\cite{CS1,Ablinger:2017err}. In the present note, we calculate the polarization function
$\hat{\tilde{\Pi}}^{(3)}(0,m_1^2,m_2^2,\mu^2)$, which is obtained as the 3rd term in the expansion
\begin{eqnarray}
\hat{\Pi}^{\mu\nu}_{ab,H}(p^2,m_1^2,m_2^2,\mu^2,\varepsilon,\hat{a}_s) 
= i (-p^2 g^{\mu\nu} + p^\mu p^\nu) \delta_{ab} \sum_{k=1}^\infty \hat{a}_s^k 
\hat{\Pi}_{H}(p^2,m_1^2,m_2^2,\mu^2,\varepsilon)   
\end{eqnarray}
in the limit $p^2 \rightarrow 0$. Here the index $H$ labels the heavy quark part of the polarization function,
$\hat{a}_s = g_s^2/(4\pi)^2$ denotes the unrenormalized strong coupling constant, 
$m_{1,2}$ are the bare heavy quark masses, and $\varepsilon = D - 4$ is the dimensional parameter. In the calculation
we refer to the Feynman rules given in \cite{YND}. 
The corresponding (single mass) expressions 
$\hat{{\Pi}}^{(1,2)}_H(0,m_1^2,m_2^2,\mu^2)$ were calculated in \cite{Kallen:1955fb,Broadhurst:1991fi,
Fleischer:1992re,Jegerlehner:1998zg,Chetyrkin:2008jk,Bierenbaum:2009mv} for QED and/or QCD.

There are six physical topologies contributing to  $\hat{\tilde{\Pi}}^{(3)}(0,m_1^2,m_2^2,\mu^2)$, Eq.~(\ref{eq:pi2m}).
The corresponding 3-loop Feynman diagrams have been generated using {\tt QGRAF} \cite{Nogueira:1991ex}
and the color-traces were calculated using {\tt Color} \cite{vanRitbergen:1998pn}. Standard Feynman 
parameter integration has been applied, cf.~\cite{Ablinger:2015tua}, representing one of the 
integrals using the Mellin-Barnes contour integral \cite{MB1a,MB1b,MB2,MB3,MB4}, also using the package
\cite{Czakon:2005rk}. The sums over the residues have been performed analytically 
using the packages {\tt Sigma} \cite{SIG1,SIG2}, {\tt EvaluateMultiSums}
and {\tt SumProduction} \cite{EMSSP}, applying procedures of {\tt HarmonicSums} 
\cite{HARMONICSUMS,Ablinger:PhDThesis,Ablinger:2011te,Ablinger:2013cf,Ablinger:2014bra}  also for the limiting 
processes in case of infinite sums.

The equal mass contributions have been calculated in 
\cite{Bierenbaum:2009mv}, Eq.~(4.7), before in case of a general 
$R_\xi$ gauge using {\tt MATAD} \cite{Steinhauser:2000ry}. The 2-mass term is given by 
\begin{eqnarray}
\hat{\tilde{\Pi}}^{(3)}(0,m_1^2,m_2^2,\mu^2) &=& \lim_{p^2\rightarrow 0}
\hat{\Pi}_{H}(p^2,m_1^2,m_2^2,\mu^2,\varepsilon)
\nonumber\\ &=& 
-C_F T_F^2 \Biggl\{
\frac{256}{9 \varepsilon^2}
+\frac{64}{3 \varepsilon} \left[\lnMa{}+\lnMb{}+\frac{5}{9}\right]
-5 \eta -\frac{5}{\eta}
\nonumber \\ &&
+\left(-\frac{5 \eta }{8}-\frac{5}{8 \eta }+\frac{51}{4}\right) \ln ^2(\eta )+\left(\frac{5}{2 \eta }-\frac{5 \eta }{2}\right) \ln (\eta)
+\frac{32 \zeta_2}{3}
\nonumber \\ &&
+32 \lnMa{} \lnMb{}
+\frac{80}{9} \lnMa{}+\frac{80}{9} \lnMb{}
+\frac{1246}{81}
\nonumber \\ &&
+\left(\frac{5 \eta^{3/2}}{2}
+\frac{5}{2 \eta^{3/2}}
+\frac{3 \sqrt{\eta}}{2}
+\frac{3}{2 \sqrt{\eta}}\right) \Biggl[\frac{1}{8} \ln \left(\frac{1+\sqrt{\eta}}{1-\sqrt{\eta}}\right) \ln^2(\eta)
\nonumber \\ &&
-{\rm Li}_3\left(-\sqrt{\eta}\right)
+{\rm Li}_3\left(\sqrt{\eta}\right)
-\frac{1}{2} \ln (\eta) \left({\rm Li}_2\left(\sqrt{\eta}\right)
-{\rm Li}_2\left(-\sqrt{\eta}\right)\right)\Biggr]
\Biggr\}
\nonumber \\ &&
-C_A T_F^2 \Biggl\{
\frac{64}{9 \varepsilon^3} \left(1+2 \xi\right)
+\frac{16}{3 \varepsilon^2} \Biggl[(1+2 \xi) \left(\lnMa{}+\lnMb{}\right)
\nonumber \\ &&
-\frac{35}{9}\Biggr]
+\frac{4}{\varepsilon} \Biggl[
\lnMa{2}+\lnMb{2}
-\frac{35}{9} \lnMa{}
-\frac{35}{9} \lnMb{}
\nonumber \\ &&
+\frac{2}{3} \zeta_2
+\frac{37}{27}
+\xi \left(\frac{4}{3} \ln^2(\eta)+4 \lnMa{} \lnMb{}+\frac{4}{3} \zeta_2+\frac{292}{81}\right)
\Biggr]
\nonumber \\ &&
+2 \left(1+\xi\right) \left(\lnMa{3}+\lnMb{3}\right)
-\frac{70}{3} \lnMa{} \lnMb{}
\nonumber \\ &&
+2 \xi \lnMa{2} \lnMb{}
+2 \xi \lnMa{} \lnMb{2}
-\frac{2}{9} \left(2+\xi\right) \ln^3(\eta)
\nonumber \\ &&
+\left(2 \left(1+2 \xi\right) \zeta_2+\frac{37}{9}+\frac{292}{27} \xi\right) \left(\lnMa{}+\lnMb{}\right)
\nonumber \\ &&
+\left[\frac{4}{3} \left(2+\xi\right) \ln(1-\eta)-\left(\eta+\frac{1}{\eta}\right) \left(\frac{2}{3}+\frac{5 \xi}{24}\right)-\frac{179}{18}-\frac{43}{36} \xi\right] \ln^2(\eta)
\nonumber \\ &&
-\frac{1}{3} \left(16+5 \xi\right) \left(\eta+\frac{1}{\eta}\right)
-\frac{70}{9} \zeta_2
-\frac{8}{9} \zeta_3 \left(7+2 \xi\right)
-\frac{3769}{243}
+\frac{262}{243} \xi
\nonumber \\ &&
+\left(\frac{1}{\eta}-\eta\right) \left(\frac{8}{3}+\frac{5}{6} \xi\right) \ln(\eta)
+\frac{8}{3} \left(2+\xi\right) \big({\rm Li}_2(\eta) \ln(\eta)-{\rm Li}_3(\eta)\big)
\nonumber \\ &&
+\left[\left(8+\frac{5}{2} \xi\right) \frac{1+\eta^3}{3 \eta^{3/2}}+\left(10+\frac{9}{2} \xi\right) \frac{1+\eta}{\sqrt{\eta}}\right]
\Biggl[\frac{1}{8} \ln \left(\frac{1+\sqrt{\eta}}{1-\sqrt{\eta }}\right) \ln^2(\eta)
\nonumber \\ &&
-{\rm Li}_3\left(-\sqrt{\eta}\right)
+{\rm Li}_3\left(\sqrt{\eta}\right)
-\frac{1}{2} \ln(\eta) \left({\rm Li}_2\left(\sqrt{\eta}\right)-{\rm Li}_2\left(-\sqrt{\eta}\right)\right)\Biggr]
\Biggr\}.
\label{eq:pi2m}
\end{eqnarray}
Here $C_A = N_c, C_F = (N_c^2-1)/(2 N_c), T_F = 1/2$
are the color factors and $N_c = 3$ in case of QCD.

Using {\tt Q2E}/{\tt Exp} \cite{Harlander:1997zb,Seidensticker:1999bb} the first terms of the 
$\eta$-expansion of (\ref{eq:pi2m}) have been obtained in Ref.~\cite{Ablinger:2017err} before. The corresponding 
expansion of (\ref{eq:pi2m}) agrees with this expression. Eq.~(\ref{eq:pi2m}) is symmetric under the
interchange $m_1 \Leftrightarrow m_2$ and depends on the gauge parameter $\xi$ at most linearly. Although
not explicitly visible in the representation given above, one can show that 
$\hat{\tilde{\Pi}}^{(3)}(0,m_1^2,m_2^2,\mu^2)$ depends only on $\eta$ and not on $\sqrt{\eta}$.

The corresponding expression in case of Quantum Electrodynamics is found by setting $C_F = 1, C_A = 0$ and
$T_F = 1$. From \cite{Lee:2013sx} the 3-loop QED result can be inferred in principle.

\noindent
{\bf Acknowledgment.}~
We would like to thank P.~Marquard for discussions and M.~Steinhauser for providing the codes {\tt MATAD 3.0} and 
{\tt Q2E}/{\tt Exp}. Discussions with A.~Behring are gratefully acknowledged. This work was supported in part by the European 
Commission through PITN-GA-2012-316704 ({HIGGSTOOLS}).

{\small
}
\end{document}